# Sharp decay estimates and numerical analysis for weakly coupled systems of two subdiffusion equations

### v1

@@ 1. Introduction and the main result
$&\partial_{t}^{\alpha}(u-u_{0})-\mathrm{div}(\bm{A}(\bm{x})\nabla u)+c_{11}(\bm{x},t)u+c_{12}(\bm{x},t)v=0,\\ &\partial_{t}^{\beta}(v-v_{0})-\mathrm{div}(\bm{B}(\bm{x})\nabla v)+c_{21}(\bm{x},t)u+c_{22}(\bm{x},t)v=0\end{aligned}&\mbox{in }\Omega\times\mathbb{R}_{+},\\ u=v=0&\mbox{on }\partial\Omega\times\mathbb{R}_{+}.\end{cases}$ (1.1)
@@
belongs to some fractional Sobolev space which only allows a pointwise definition for $\alpha>1/2$. Then the notation $\partial_{t}^{\alpha}(u-u_{0})$ means $u(\bm{x},\,\cdot\,)-u_{0}(\bm{x})\in H_{\alpha}(0,T)$ for a.e. $\bm{x}\in\Omega$ and any $T>0$, and the initial condition $u=u_{0}$ in $\Omega\times\{0\}$ only makes usual sense for $\alpha>1/2$. Especially, for $\alpha=1$ one can interpret $\partial_{t}^{\alpha}(u-u_{0})=\partial_{t}u$ as the usual partial derivative in time. On the spatial direction of (1.1), we assume that the principal coefficients $\bm{A},\bm{B}$ in the elliptic parts and the zeroth order coefficients $c_{k\ell}$ ($k,\ell=1,2$) satisfy
@@
The physical background of fractional diffusion equations comes from the extension of classical diffusion models. The introduction of fractional derivatives effectively captures anomalous phenomenon such as non-Gaussian profiles and long-term memory effects in porous media and biological tissues, which cannot be well described by traditional integer-order diffusion equations, see e.g., [2], [5] and the references therein. Fractional diffusion systems employ nonlocal operators and coupling mechanisms to not only describe fundamental diffusion processes such as single equations but also incorporate complex factors, which have significant applications in fields such as chemical reactions, biological mass transfer, and environmental engineering. For example, in the transport of solute through porous media, they can integrate multiple influencing factors into a unified model and, therefore, can describe the dynamic mass exchange process between mobile and immobile zones [3], [12], [18], [19] and the references therein.
The long-time asymptotic behavior is the most remarkable difference between fractional and non-fractional equations. This fundamental difference was first quantitatively characterized through single time-fractional equation studies: Sakamoto and Yamamoto [17] demonstrated that solutions to time-fractional diffusion equations exhibit fractional polynomial decay, where the memory effect fundamentally alters the exponential decay pattern of classical diffusion. Subsequent extensions to multi-term fractional models revealed that the lowest order dominates the long-term decay [11]. Remarkably, Luchko et al. [14] and Kubica and Ryszewska [7] observed logarithmic-type decay in distributed-order models, which generalize the multi-term time-fractional diffusion equation (featuring a finite sum of fractional derivatives) to a framework involving continuous integration over infinitely many fractional derivatives of distinct orders. Vergara and Zacher [20] established the quantitative relationship between decay rates and fractional order in single equations with time-dependent coefficients. Li, Huang, Liu [9] studied the same topic for a weakly coupled subdiffusion systems of $K$ components
@@
where $01>\alpha_{1}\geq\cdots\geq\alpha_{K}>$ 0, $\bm{A}_{k}$ satisfy the same assumption as (1.2) and $c_{k\ell}\in L^{\infty}(\Omega)$ depend only on $\bm{x}$ ($k,\ell=1,\dots,K$). Under a certain non-positivity assumption on $c_{k\ell}$, [10, Theorem 2] established the decay estimate (see also Lemma 2.5)
@@
Keeping the above background in mind, in this paper we are interested in sharpening the decay estimate (1.5) under the aforementioned situation. As the first attempt, we restrict ourselves to the formulation (1.1) of two components. Since $v_{0}\equiv 0$ in $\Omega$, the time evolution of $v$ relies completely on the supply from $u$. Therefore, one can imagine that the higher order $\alpha$ may appear to influence the long-time behavior of the solution. To verify our conjecture, we perform numerical tests under various settings, including both cases of $\alpha<1$ and $\alpha=1$. As is reported in Section 5, we observed a decay rate of $t^{-\alpha}$ for $\alpha<1$ as predicted. Surprisingly, we witnessed an unexpected decay pattern for $\alpha=1$ which seems never be discovered in literature. These observations motivate us to establish the following main result of this manuscript.
@@
Then there exists a constant $C>0$ such that the solution $(u,v)$ to (1.1) satisfy
@@
The remainder of the paper is organized as follows. In Section 2, several necessary results are given, including asymptotic estimates of Mittag-Leffler functions, the coercivity of fractional derivatives, and the residue theory on the complex plane cutting off the negative real axis. In Section 3, we establish the maximum principle and the asymptotic behavior of the solution to a coupled fractional ordinary system, from which we further employ the coercivity to give a proof of the main theorem. In Section 4, we propose semi-implicit and fully implicit numerical schemes for fractional diffusion equations and perform stability analysis to ensure the reliability of numerical simulations. Finally, in Section 5, we verify the validity of the theoretical results through specific numerical experiments demonstrating the consistency between the numerical results and the theoretical analysis.
@@ 2. Preliminaries
[…]
We first recall the frequently used Mittag-Leffler functions $E_{\eta,\mu}(z)$ of two parameters $\eta>0$, $\mu\in\mathbb{R}$ defined by (e.g. Podlubny [16])
@@
Many properties of the Mittag-Leffler functions play important roles in fractional calculus, and here we only invoke their estimate and positivity for our use in this article.
@@
    Let $20<\eta<$ 2 and $\mu\in\mathbb{R}$. Then there exists a constant $C>0$ depending only on $\eta,\mu$ such that
@@
    $0<E_{\eta,1}(-z)\leq 1,\quad 0<E_{\eta,\eta}(-z)\leq 1,\quad\forall\,z\geq 0$
@@
Next, we turn to the coercivity of the fractional derivative $\partial_{t}^{\alpha}$. For the usual first order partial derivative $\partial_{t}$, it is readily seen for a smooth function $w$ define in $\Omega\times\mathbb{R}_{+}$ that
@@
We shall prove the above inequality by a density argument. To this end, we first assume $w\in C^{1}([0,T];L^{2}(\Omega))$. Then it follows from Li, Huang and Yamamoto [10, Lemma 4.1] that
@@
$w_{m}-w_{0}\in{}_{0}C^{1}([0,T];L^{2}(\Omega)),\quad w_{m}-w_{0}\longrightarrow w-w_{0}\mbox{ in }H_{\alpha}(0,T;L^{2}(\Omega))\mbox{ as }m\to\infty.$
@@
Then we can conclude (2.1) by simply passing $m\to\infty$ in the above inequality. ∎
@@
Now for a multi-valued meromorphic function $g$ satisfying $|g(z)|\leq C|z|^{-p}$ with some constant $p>0$, suppose that all poles $\{z_{1},\dots,z_{N}\}$ of $g(z)$ are distributed in the left half complex plane excluding the complex real axis. On the basis of Lemma 2.3, we establish a more general residue result which is suitable also for the function $g(z)$ on the complex plane cutting off the negative axis.
@@
Let $g$ be a meromorphic function on the complex plane cutting off the negative axis. Assume that there exist a constant $p>0$ such that
@@
Finally, passing $\varepsilon\to 0$, we can conclude the desired residue theorem on the complex plane cutting off the negative axis, which finishes the proof of Lemma 2.4. ∎
@@
In [9], all fractional orders of time derivatives were assumed to be strictly smaller than $1$. In this sense, Lemma 2.5 fails to cover the special case of $\alpha=1$ in (1.1). However, the argument of an iterative construction of the mild solution definitely works for $\alpha=1$ and thus the well-posedness part in Lemma 2.5(i) still holds. On the other hand,
@@ 3.1. Maximum principle of fractional ordinary systems
$\begin{cases}\partial_{t}^{\alpha}(U-a)+\eta_{1}U-\mu_{1}V=F,\\ \partial_{t}^{\beta}(V-b)-\mu_{2}U+\eta_{2}V=G\end{cases}\mbox{in }\mathbb{R}_{+},$ (3.1)
@@
Assume $\mu_{2}>0,$ $a,b,\eta_{1},\eta_{2},\mu_{1},\mu_{2}\geq 0$ and $0\leq F,G\in L_{\mathrm{loc}}^{2}(\overline{\mathbb{R}_{+}})$. Then there exists a unique solution $(U,V)\in L_{\rm loc}^{2}(\overline{\mathbb{R}_{+}})$ to (3.1) such that $U\geq 0,V\geq 0$ in $\overline{\mathbb{R}_{+}}$. Especially, if we further assume $a>0,$ then $U>0,V>0$ in $\mathbb{R}_{+}$.
@@
$\begin{cases}\partial_{t}^{\alpha}(U-a)+\eta_{1}U=F+\mu_{1}V,\\ \partial_{t}^{\beta}(V-b)+\eta_{2}V=G+\mu_{2}U\end{cases}\mbox{in }\mathbb{R}_{+}.$
Employing the Mittag-Leffler function, we can easily see that $(U,V)$ satisfies the integral equation
@@
[…]
@@
[…]
@@
Therefore, the equation (3.4) for $U_{m},V_{m}$, $m=0,1,\cdots$, can be rephrased by
@@
To this end, we shall give some estimates for $\bm{U}_{m}$ in some appropriate function space. Letting $T>0$ be fixed arbitrarily, by means of Lemma 2.1, we assert that $\bm{U}_{m}\in L^{2}(0,T)$ for $m=0,1,\dots$ and there exists a constant $C_{K}$ such that
@@
where the constant $C_{K}>0$ depends on $\eta_{1},\eta_{2},\mu_{1},\mu_{2},\alpha,\beta$ and $T$. Now we note the inductive assumption to derive
@@
in view of the semi-group property $J^{\gamma_{1}}J^{\gamma_{2}}=J^{\gamma_{1}+\gamma_{2}}$ for any $\gamma_{1},\gamma_{2}>0$. By the induction argument, we see that the estimate (3.5) holds for any $m=1,2,\dots$. Again, from Young’s convolution inequality, we see that $\bm{U}_{m}=\sum_{j=0}^{m}(\bm{U}_{j+1}-\bm{U}_{j})\in L^{2}(0,T)$ for any $m=1,2,\dots$.
@@
Next, we show that $\bm{U}$ must be strictly positive in $(0,T]$ if $a>0$. For this, we note that $\bm{U}$ satisfies the integral equation
@@
Moreover, from the facts in Lemma 2.1, it follows that $a\,E_{\alpha,1}(-\eta_{1}t^{\alpha})>0$. Collecting all the above results, we finally see that $\bm{U}(t)>0$ for any $t>0$. We finish the proof of the lemma. ∎
@@ 3.2. Asymptotic behavior of fractional ordinary system
$\begin{cases}\partial_{t}^{\alpha}(U-1)+c_{1}U-c_{2}V=0,\\ \partial_{t}^{\beta}V-c_{2}U+c_{1}V=0\end{cases}\mbox{in }\mathbb{R}_{+},$ (3.6)
@@
$10<U(t)+V(t)\leq\begin{cases}C\,t^{-\alpha},&\alpha<1,\\ C\,t^{-(1+\beta)},&\alpha=1,\end{cases}\quad\forall\,t>$ 1 .
@@
Taking the Laplace transform $\widehat{f}(s)=\int_{\mathbb{R}_{+}}\mathrm{e}^{-st}f(t)\,\mathrm{d}t$ on both sides of (3.6) yileds
@@
It is obvious that $I_{1}(t)$ and $I_{3}(t)$ are of exponential decay as $t\to+\infty$. For $I_{2}(t)$ and $I_{4}(t)$, we have
@@
To deal with the integrand of $I_{4}(t)$, we substitute (3.7) into $\mathrm{Im}[\widehat{V}(s)]_{s=r\,\mathrm{e}^{\mathrm{i}\,\pi}}$ to deduce
@@
$q(r)$ =$=\{(c_{1}+r^{\alpha}\cos\alpha\pi)+\mathrm{i}\,r^{\alpha}\sin\alpha\pi\}\{(c_{1}+r^{\beta}\cos\beta\pi)+\mathrm{i}\,r^{\beta}\sin\beta\pi\}-c_{2}^{2}$
=$=\left\{(c_{1}^{2}-c_{2}^{2})+c_{1}(r^{\alpha}\cos\alpha\pi+r^{\beta}\cos\beta\pi)+r^{\alpha+\beta}\cos(\alpha+\beta)\pi\right\}$
@@
In fact, Since $\widehat{V}(s)$ has no pole on the negative real axis, it is impossible for $q(r)$ to have zero points in $\mathbb{R}_{+}$. For $r=0$, it follows from the assumption $c_{1}>c_{2}>0$ that $|q(0)|^{2}=(c_{1}^{2}-c_{2}^{2})^{2}>0$. Meanwhile, we readily see from the above expression (3.11) of $q(r)$ that
@@
where we employed Gamma function to calculate e.g.
@@
On the other hand, the calculation of $I_{2}(t)$ is more complicated but the methodology is identical. We substitute (3.7) into $\mathrm{Im}[\widehat{U}(s)]_{s=r\,\mathrm{e}^{\mathrm{i}\,\pi}}$ to deduce
@@
=$=\left\{c_{1}\cos\alpha\pi+r^{\beta}\cos(\alpha+\beta)\pi\right\}+\mathrm{i}\left\{c_{1}\sin\alpha\pi+r^{\beta}\sin(\alpha+\beta)\pi\right\}.$
@@
=$=\left\{c_{1}\sin\alpha\pi+r^{\beta}\sin(\alpha+\beta)\pi\right\}\left\{(c_{1}^{2}-c_{2}^{2})+c_{1}(r^{\alpha}\cos\alpha\pi+r^{\beta}\cos\beta\pi)+r^{\alpha+\beta}\cos(\alpha+\beta)\pi\right\}$
@@
Then we combine the above inequality with (3.12) and (3.14) to estimate $I_{2}(t)$ in (3.8) as
@@
00 0 $=(u(t),\partial_{t}^{\alpha}(u(t)-u_{0})-\mathrm{div}(\bm{A}\nabla u(t))+c_{11}(t)u(t)+c_{12}(t)v(t))$
@@
where $C_{\Omega}$ and $\kappa_{1}$ were the constants introduced in (1.6) and (1.7), respectively. By $U(t)\geq 0$, we have
$\partial_{t}^{\alpha}(U-U(0))+\frac{\kappa_{0}}{C_{\Omega}^{2}}U-\kappa_{1}V\leq 0\quad\mbox{in }\mathbb{R}_{+}.$ (3.15)
Repeating the same procedure for the governing equation of $v$ in (1.1), we obtain
$\partial_{t}^{\beta}V+\frac{\kappa_{0}}{C_{\Omega}^{2}}V-\kappa_{1}U\leq 0\quad\mbox{in }\mathbb{R}_{+},$ (3.16)
@@
As $(U,V)$ satisfies the coupled system (3.15)–(3.16) of fractional ordinary differential inequalities, for clarity we introduce an auxiliary fractional ordinary differential system
$&\partial_{t}^{\alpha}(\overline{U}-1)+\frac{\kappa_{0}}{C_{\Omega}^{2}}\overline{U}-\kappa_{1}\overline{V}=0,\\ &\partial_{t}^{\alpha}\overline{V}+\frac{\kappa_{0}}{C_{\Omega}^{2}}\overline{V}-\kappa_{1}\overline{U}=0\end{aligned}\quad\mbox{in }\mathbb{R}_{+}.\right.$
@@
$U\leq U(0)\overline{U},\quad V\leq U(0)\overline{V}\quad\mbox{in }\mathbb{R}_{+}.$
On the other hand, owing to (1.7), the key assumption in Lemma 3.2 is satisfied, from which we conclude

### v2

@@ 1. Introduction and the main result
$\begin{cases}\begin{aligned} &\partial_{t}^{\alpha}(u-u_{0})-\mathrm{div}(\bm{A}(\bm{x})\nabla u)+c_{11}(\bm{x},t)u+c_{12}(\bm{x},t)v=0,\\ &\partial_{t}^{\beta}(v-v_{0})-\mathrm{div}(\bm{B}(\bm{x})\nabla v)+c_{21}(\bm{x},t)u+c_{22}(\bm{x},t)v=0\end{aligned}&\mbox{in }\Omega\times\mathbb{R}_{+},\\ u=v=0&\mbox{on }\partial\Omega\times\mathbb{R}_{+}.\end{cases}$ (1.1)
@@
belongs to some fractional Sobolev space which only allows a pointwise definition for $\alpha>1/2$. Then the notation $\partial_{t}^{\alpha}(u-u_{0})$ means $u(\bm{x},\,\cdot\,)-u_{0}(\bm{x})\in H_{\alpha}(0,T)$ for a.e. $\bm{x}\in\Omega$ and any $T>0$, and the initial condition $u=u_{0}$ in $\Omega\times\{0\}$ only makes usual sense for $\alpha>1/2$. Especially, for $\alpha=1$ one can interpret $\partial_{t}^{\alpha}(u-u_{0})=\partial_{t}u$ as the usual partial derivative in time. In the spatial direction of (1.1), let the principal coefficients $\bm{A},\bm{B}$ in the elliptic parts and the zeroth order coefficients $c_{k\ell}$ ($k,\ell=1,2$) satisfy
@@
The physical background of fractional diffusion equations comes from the extension of classical diffusion models. The introduction of fractional derivatives effectively captures anomalous phenomenon such as non-Gaussian profiles and long-term memory effects in porous media and biological tissues, which cannot be well described by traditional integer-order diffusion equations, see e.g., [2, 5, 16] and the references therein. Fractional diffusion systems employ nonlocal operators and coupling mechanisms to not only describe fundamental diffusion processes such as single equations but also incorporate complex factors, which have significant applications in fields such as chemical reactions, biological mass transfer, and environmental engineering. For example, in the transport of solute through porous media, they can integrate multiple influencing factors into a unified model and, therefore, can describe the dynamic mass exchange process between mobile and immobile zones [3, 12, 19, 20] and the references therein.
The long-time asymptotic behavior is the most remarkable difference between fractional and non-fractional equations. This fundamental difference was first quantitatively characterized through single time-fractional equation studies: Sakamoto and Yamamoto [18] demonstrated that solutions to time-fractional diffusion equations exhibit fractional polynomial decay, where the memory effect fundamentally alters the exponential decay pattern of classical diffusion. Subsequent extensions to multi-term fractional models in [11] revealed that the lowest order dominates the long-term decay. Remarkably, Luchko et al. [14] and Kubica and Ryszewska [7] observed logarithmic-type decay in distributed-order models, which generalize the multi-term time-fractional diffusion equation (featuring a finite sum of fractional derivatives) to a framework involving continuous integration over infinitely many fractional derivatives of distinct orders. Vergara and Zacher [21] established the quantitative relation between decay rates and fractional orders in single equations with time-dependent coefficients. Li, Huang, Liu [9] studied the same topic for a weakly coupled subdiffusion systems of $K$ components
@@
where $1>\alpha_{1}\geq\cdots\geq\alpha_{K}>0$, $\bm{A}_{k}$ satisfy the same assumption as (1.2) and $c_{k\ell}\in L^{\infty}(\Omega)$ depend only on $\bm{x}$ ($k,\ell=1,\dots,K$). Under a certain non-positivity assumption on $c_{k\ell}$, [9, Theorem 2] established the decay estimate (see also Lemma 2.5)
@@
Keeping the above background in mind, in this paper we are interested in sharpening the decay estimate (1.5) under the aforementioned situation. As the first attempt, we restrict ourselves to the formulation (1.1) of two components. Since $v_{0}\equiv 0$ in $\Omega$, the time evolution of $v$ relies completely on the supply from $u$. Therefore, one can imagine that the higher order $\alpha$ may appear to influence the long-time behavior of the solution. To verify our conjecture, we perform numerical tests under various settings, including both cases of $\alpha<1$ and $\alpha=1$. As is reported in Section 5, we observed a decay rate of $t^{-\alpha}$ for $\alpha<1$ as predicted. Surprisingly, we witnessed an unexpected decay pattern for $\alpha=1$ which seems to be never discovered in the literature. These observations motivate us to establish the following main result of this manuscript.
@@
Then there exists a constant $C>0$ such that the solution $(u,v)$ to (1.1) satisfies
@@
The remainder of the paper is organized as follows. In Section 2, several necessary results are given, including asymptotic estimates of Mittag-Leffler functions, the coercivity of fractional derivatives, and the residue theory on the complex plane cutting off the negative real axis. In Section 3, we establish the maximum principle and the asymptotic behavior of the solution to a coupled fractional ordinary system, from which we further employ the coercivity to give a proof of the main theorem. In Section 4, we propose semi-implicit and fully implicit numerical schemes for fractional diffusion equations, and perform stability analysis to ensure the reliability of numerical simulations. Finally, in Section 5, we verify the validity of the theoretical results through specific numerical experiments demonstrating the consistency between the numerical results and the theoretical analysis.
@@ 2. Preliminaries
In the sequel, by $\|\cdot\|$ and (,)($\cdot\,$$\cdot\,$) we denote the norm and the inner product of $L^{2}(\Omega)$ respectively. In addition, by $C>0$ we denote generic constants which may change from line to line.
We first recall the frequently used Mittag-Leffler functions $E_{\eta,\mu}(z)$ with two parameters $\eta>0$, $\mu\in\mathbb{R}$ defined by (e.g. Podlubny [17])
@@
Many properties of the Mittag-Leffler functions play important roles in fractional differential equations, and here we only invoke their estimates and positivity for our use in this article.
@@
    Let $0<\eta<2$ and $\mu\in\mathbb{R}$. Then there exists a constant $C>0$ depending only on $\eta,\mu$ such that
@@
    $0<E_{\eta,1}(-z)\leq 1,\quad 0<E_{\eta,\eta}(-z)\leq 1,\quad\forall\,z\geq 0.$
@@
Next, let us turn to the coercivity of the fractional derivative $\partial_{t}^{\alpha}$. For the usual first order partial derivative $\partial_{t}$, it is readily seen for a smooth function $w$ defined in $\Omega\times\mathbb{R}_{+}$ that
@@
The above inequality can be proved by a density argument. To this end, we first assume $w\in C^{1}([0,T];L^{2}(\Omega))$. Then it follows from Li, Huang and Yamamoto [10, Lemma 4.1] that
@@
[…]
@@
Then the coercivity (2.1) can be derived by simply passing $m\to\infty$ in the above inequality. ∎
@@
Now for a multi-valued meromorphic function $g$ satisfying $|g(z)|\leq C|z|^{-p}$ with some constant $p>0$, suppose that all poles $\{z_{1},\dots,z_{N}\}$ of $g(z)$ are distributed in the left half complex plane excluding the real axis. On the basis of Lemma 2.3, we establish a more general residue result which is suitable also for the function $g(z)$ on the complex plane cutting off the negative axis.
@@
Let $g$ be a meromorphic function on the complex plane cutting off the negative axis. Assume that there exists a constant $p>0$ such that
@@
Finally, passing $\varepsilon\to 0$, it arrives at the desired residue theorem on the complex plane cutting off the negative axis, which finishes the proof of Lemma 2.4. ∎
@@
In [9], all fractional orders of time derivatives were assumed to be strictly smaller than $1$. In this sense, Lemma 2.5 fails to cover the special case of $\alpha=1$ in (1.1). However, the argument of an iterative construction of the mild solution definitely works for $\alpha=1$ and thus the well-posedness part in Lemma 2.5(i) still holds.
@@ 3.1. Maximum principle of fractional ordinary systems
[…]
@@
Assume $\mu_{2}>0,$ $a,b,\eta_{1},\eta_{2},\mu_{1},\mu_{2}\geq 0$ and $0\leq F,G\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. Then there exists a unique solution $(U,V)\in L_{\rm loc}^{2}(\mathbb{R}_{+})$ to (3.1) such that $U\geq 0,V\geq 0$ in $\mathbb{R}_{+}$. Especially, if we further assume $a>0,$ then $U>0,V>0$ in $\mathbb{R}_{+}$.
@@
Employing the Mittag-Leffler function, it can be easily seen that $(U,V)$ satisfies the integral equation
$\left\{\begin{aligned} U(t)&=U_{1}(t)+\mu_{1}\int_{0}^{t}\tau^{\alpha-1}E_{\alpha,\alpha}(-\eta_{1}\tau^{\alpha})V(t-\tau)\,\mathrm{d}\tau,\\ V(t)&=V_{1}(t)+\mu_{2}\int_{0}^{t}\tau^{\beta-1}E_{\beta,\beta}(-\eta_{2}\tau^{\beta})U(t-\tau)\,\mathrm{d}\tau,\end{aligned}\right$. (3.2)
@@
$\left\{\begin{aligned} U_{1}(t)&:=a\,E_{\alpha,1}(-\eta_{1}t^{\alpha})+\mu_{1}\int_{0}^{t}\tau^{\alpha-1}E_{\alpha,\alpha}(-\eta_{1}\tau^{\alpha})F(t-\tau)\,\mathrm{d}\tau,\\ V_{1}(t)&:=b\,E_{\beta,1}(-\eta_{2}t^{\beta})+\mu_{2}\int_{0}^{t}\tau^{\beta-1}E_{\beta,\beta}(-\eta_{2}\tau^{\beta})G(t-\tau)\,\mathrm{d}\tau.\end{aligned}\right$. (3.3)
@@
$\left\{\begin{aligned} U_{m+1}(t)&=U_{1}(t)+\mu_{1}\int_{0}^{t}\tau^{\alpha-1}E_{\alpha,\alpha}(-\eta_{1}\tau^{\alpha})V_{m}(t-\tau)\,\mathrm{d}\tau,\\ V_{m+1}(t)&=V_{1}(t)+\mu_{2}\int_{0}^{t}\tau^{\beta-1}E_{\beta,\beta}(-\eta_{2}\tau^{\beta})U_{m}(t-\tau)\,\mathrm{d}\tau,\end{aligned}\right.\quad m=0,1,\dots.$ (3.4)
@@
Therefore, (3.4) can be rephrased as
@@
Next, we shall give some estimates for $\bm{U}_{m}$ in some appropriate function space. Letting $T>0$ be fixed arbitrarily, by means of Lemma 2.1, we assert that $\bm{U}_{m}\in L^{2}(0,T)$ for $m=0,1,\dots$ and there exists a constant $C_{K}$ such that
@@
where the constant $C_{K}>0$ depends on $\eta_{1},\eta_{2},\mu_{1},\mu_{2},\alpha,\beta$ and $T$. Now we use the inductive assumption to derive
@@
in view of the semi-group property $J^{\gamma_{1}}J^{\gamma_{2}}=J^{\gamma_{1}+\gamma_{2}}$ for any $\gamma_{1},\gamma_{2}>0$. By the induction argument, we see that the estimate (3.5) holds for any $m=1,2,\dots$. Again, from Young’s convolution inequality, it follows that $\bm{U}_{m}=\sum_{j=0}^{m}(\bm{U}_{j+1}-\bm{U}_{j})\in L^{2}(0,T)$ for any $m=1,2,\dots$.
@@
Finally, we show that $\bm{U}$ must be strictly positive in $(0,T)$ if $a>0$. For this, we note that $\bm{U}$ satisfies the integral equation
@@
Moreover, from the facts in Lemma 2.1, it follows that $a\,E_{\alpha,1}(-\eta_{1}t^{\alpha})>0$. Collecting all the above results, we conclude that $\bm{U}(t)>\bm{0}$ for any $t>0$. The proof of the lemma is complete. ∎
@@ 3.2. Asymptotic behavior of fractional ordinary system
[…]
@@
$0<U(t)+V(t)\leq\begin{cases}C\,t^{-\alpha},&\alpha<1,\\ C\,t^{-(1+\beta)},&\alpha=1,\end{cases}\quad\forall\,t>1.$
@@
Taking the Laplace transform $\widehat{f}(s)=\int_{\mathbb{R}_{+}}\mathrm{e}^{-st}f(t)\,\mathrm{d}t$ on both sides of (3.6) yields
@@
It is obvious that $I_{1}(t)$ and $I_{3}(t)$ are of exponential decay as $t\to+\infty$. For $I_{2}(t)$ and $I_{4}(t)$, it can be easily proved that
@@
To deal with the integrand of $I_{4}(t)$, the substitution of (3.7) into $\mathrm{Im}[\widehat{V}(s)]_{s=r\,\mathrm{e}^{\mathrm{i}\,\pi}}$ results in
@@
$q(r)$ $=\{(c_{1}+r^{\alpha}\cos\alpha\pi)+\mathrm{i}\,r^{\alpha}\sin\alpha\pi\}\{(c_{1}+r^{\beta}\cos\beta\pi)+\mathrm{i}\,r^{\beta}\sin\beta\pi\}-c_{2}^{2}$
$=\left\{(c_{1}^{2}-c_{2}^{2})+c_{1}(r^{\alpha}\cos\alpha\pi+r^{\beta}\cos\beta\pi)+r^{\alpha+\beta}\cos(\alpha+\beta)\pi\right\}$
@@
In fact, since $\widehat{V}(s)$ has no pole on the negative real axis, it is impossible for $q(r)$ to have zero points in $\mathbb{R}_{+}$. For $r=0$, it follows from the assumption $c_{1}>c_{2}>0$ that $|q(0)|^{2}=(c_{1}^{2}-c_{2}^{2})^{2}>0$. Meanwhile, it is readily seen from the above expression (3.11) of $q(r)$ that
@@
where the Gamma function is employed to calculate, e.g.
@@
On the other hand, the calculation of $I_{2}(t)$ is more complicated, but the methodology is identical. Substituting (3.7) into $\mathrm{Im}[\widehat{U}(s)]_{s=r\,\mathrm{e}^{\mathrm{i}\,\pi}}$ leads to
@@
$=\left\{c_{1}\cos\alpha\pi+r^{\beta}\cos(\alpha+\beta)\pi\right\}+\mathrm{i}\left\{c_{1}\sin\alpha\pi+r^{\beta}\sin(\alpha+\beta)\pi\right\}.$
@@
$=\left\{c_{1}\sin\alpha\pi+r^{\beta}\sin(\alpha+\beta)\pi\right\}\left\{(c_{1}^{2}-c_{2}^{2})+c_{1}(r^{\alpha}\cos\alpha\pi+r^{\beta}\cos\beta\pi)+r^{\alpha+\beta}\cos(\alpha+\beta)\pi\right\}$
@@
The above inequality is then combined with (3.12) and (3.14) to estimate $I_{2}(t)$ in (3.8) as
@@
$0$ $=(u(t),\partial_{t}^{\alpha}(u(t)-u_{0})-\mathrm{div}(\bm{A}\nabla u(t))+c_{11}(t)u(t)+c_{12}(t)v(t))$
@@
where $C_{\Omega}$ and $\kappa_{1}$ are the constants introduced in (1.6) and (1.7), respectively. From $U(t)\geq 0$, it follows that
[…]
Repeating the same procedure for the governing equation of $v$ in (1.1) implies that
[…]
@@
As $(U,V)$ satisfies the coupled system (3.15)–(3.16) of fractional ordinary differential inequalities, for clarity, we introduce an auxiliary fractional ordinary differential system
[…]
@@
[…]
On the other hand, due to (1.7), the key assumption in Lemma 3.2 is satisfied, hence it can be inferred that